# PointNetPGAP-SLC: A 3D LiDAR-based Place Recognition Approach with Segment-level Consistency Training for Mobile Robots in Horticulture

### v1

@@ IV-E1 Ablation Study
The results of the ablation study are reported in Table IV. For this study, we report the Recall@1 and Recall@1% of each aggregation approach separately and combined. The results indicate that combining both aggregators yields higher performance compared to each aggregator individually, improving performance by 0.6 percentage points (pp) for retrieving the Top-1% candidates and 4.7 pp for retrieving the Top-1 candidate. These results suggest that the real impact of combining these aggregation approaches is on modeling the closest neighbor.
TABLE III: Recall@1 performance of the models trained: (without SLC)/(with SLC). The bolded scores represent the highest value between trained without and with SLC, while the underling indicates the highest score in each column.
@@
TABLE IV: Ablation study of PNPGAP, where GAP represents the global average pooling aggregator and PFI represents the aggregator based on pairwise feature interactions. The bolded scores represent the highest values in each column.
IV-E2 Comparison to SOTA
@@
PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD (the second-best model) by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the top-1 candidate. This is also evident in Fig. 6, where a qualitative assessment is shown on the paths, displaying the true positive predictions of each model when retrieving the top-1 candidate.
@@
When analyzing the performance of retrieving the top-25 candidates (Fig. 4), LOGG3D-Net, which relies on a voxel-based network for feature extraction, exhibits the highest increase in performance. Meanwhile, OverlapTransformer, which relies on range images and on a CNN-based network for feature extraction, shows poor performance across all sequences. The inadequate performance of OverlapTransformer is also apparent in Fig. 6, where in some segments the model fails to predict any loops correctly. This suggests the unsuitability of this model for agricultural environments.
At the segment-level, most models perform best in segments 3 and 4, representing the extremities of the crops. As illustrated in Fig. 1, these segments are distinctive, while segments representing the rows exhibit higher similarities and thus create ambiguities among neighboring rows, resulting in lower retrieval performance.
Figure 6: True positive predictions for retrieving the top-1 candidate.
IV-E3 Training with the SLC Model
@@
The results when adding the SLC model at training time (presented in Table III) indicate that this additional training signal improves the performance throughout all sequences and models (except OverlapTransformer), with an average performance increase of 3.7 pp for LOGG3D, 4.3 pp for PointNetVLAD, and 3.34 pp for the proposed PNPGAP model.
These findings can be further analyzed in more detail in Fig. 3, where a comparison between PNPGAP and PNPGAP-SLC is made at the segment-level. The results show that, while the performance in segments with high performance remains mostly unchanged, segments such as those corresponding to rows (i.e., 1, 2, and 5) increase performance, indicating that the SLC model produces better descriptors that are less susceptible to ambiguities.
@@ IV-F Runtime Analysis
TABLE V: The average runtime per batch, where the batch size is 20 scans.
@@
The processing capacity and model size are crucial factors for real-world implementation, particularly in resource-constrained robotics applications. Throughout the runtime experiment, the results presented in Table V were obtained using identical hardware specifications as detailed in Section IV-D: an AMD Ryzen 9 5900X 12-Core CPU with 64 GB RAM and a GeForce RTX 3090 24GB GPU. The reported outcomes specifically denote the duration, expressed in milliseconds (ms), that each model required for processing a batch of 20 scans. The retrieval duration was omitted from this experiment because it remains uniform across all models. The size of the models is reported through the count of parameters, denominated in millions (M). PNPGAP emerges as one of the smaller models, comprising 0.4 million parameters and demanding on average 4.4 ms per batch.
A comparative analysis against other SOTA models reveals that PNPGAP has approximately 5% of the size of LOGG3D-Net and is 31 times faster. Furthermore, in comparison to PointNetVLAD, PNPGAP is only 2% of its size and is 6 times faster. These outcomes highlight the computational efficiency of the proposed PNPGAP model.
@@ V CONCLUSIONS
In this letter, we addressed the 3D LiDAR place recognition problem in horticultural settings by introducing a novel aggregation approach that fuses two statistic-inspired aggregators into a model called PointNetPGAP; an SLC model to handle inter-row descriptor ambiguities; and a novel dataset, HORTO-3DLM, which includes four sequences: three from orchards and one from a strawberry plantation.
The experimental evaluation was conducted on the HORTO-3DLM dataset, and the proposed PointNetPGAP model was benchmarked against SOTA models. An ablation study demonstrated the superiority of the combined aggregators, while the results at both the sequence and segment levels show the proposed model outperforming SOTA models while being more efficient.
Regarding the proposed training approach, all models were trained with and without the SLC model. The results showed that, except for one model, incorporating the SLC model during training led to an increase in performance. This improvement was primarily observed in segments with high inter-row similarity. This suggests that providing a context-specific training signal, as offered by the SLC model, may be crucial for effectively training place recognition models in environments where LiDAR scans overlap significantly, even when originating from different rows.

### v2

@@ IV-E1 Ablation Study
The ablation study is conducted following the cross-validation protocol outlined in Section IV-C1, with results presented in Table V. For this analysis, we report Recall@1 and Recall@1% for each aggregation approach, both individually and combined. The results suggest that combining both aggregators results in superior performance compared to using each aggregator separately, with an improvement of 0.6 percentage points (pp) in retrieving the Top-1% candidates and 4.7 pp for retrieving the Top-1 candidate. These findings indicate that the proposed aggregator significantly enhances the model’s ability to identify the nearest neighbor.
@@
IV-E2 Cross-Validation (same domain)
@@
The results of this experiment are presented in Table IV. PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the Top-1 candidate.This superiority is further illustrated in Fig. 6, where a qualitative assessment shows the true positive predictions of each model for the Top-1 candidate retrieval along the paths. Given that PNPGAP and PointNetVLAD share the same backbone, these results suggest that the proposed aggregator is more effective in horticultural environments, particularly in handling sparse scans, than other first- and second-order aggregators.
@@
When analyzing the performance of retrieving the Top-25 candidates (Fig. 4), LOGG3D-Net, which relies on a voxel-based network for feature extraction, exhibits the highest increase in performance. Meanwhile, OverlapTransformer, which relies on range images and on a CNN-based network for feature extraction, shows poor performance across all sequences. The low performance of OverlapTransformer is also apparent in Fig. 6, where in some segments the model fails to predict any loops correctly. This suggests the unsuitability of this model for agricultural environments.
@@
When adding the SLC model at training time (as shown in Table IV), the results indicate that this additional training signal enhances performance across all sequences and models (except for OverlapTransformer), with an average performance increase of 3.7 pp for LOGG3D, 4.3 pp for PointNetVLAD, and 3.34 pp for the proposed PNPGAP model.
At the segment-level (see Fig.5), most models (without SLC) achieve their highest performance in segments 3 and 4, which represent the extremities of the crops. As illustrated in Fig. 1,these segments are more distinctive. In contrast, segments representing the rows exhibit greater similarity, creating ambiguities among adjacent rows, which leads to lower retrieval performance.
However, when the SLC model is incorporated, the results at the segment level indicate a notable performance increase, particularly in the segments corresponding to rows (i.e., 0, 1, 2, and 5). This improvement suggests that the SLC model produces more robust descriptors that are less susceptible to ambiguities.
Figure 6: True positive predictions of the top-1 retrieved candidates.
TABLE VI: Recall@1 of the cross-domain evaluation: (without SLC)/(with SLC). The bolded scores represent the highest value between trained without and with SLC, while the underlined scores indicate the highest score in each column.
  Validation Testing 00 02 05 06 08 MEAN Overlap Transformer [5] 0.247/0.278+0.031 0.192/0.253+0.061 0.264/0.361+0.097 0.227/0.355+0.128 0.023/0.040+0.017 0.176/0.252+0.076 LOGG3D [7] 0.043/0.083+0.040 0.084/0.116+0.032 0.108/0.176+0.068 0.271/0.209-0.062 0.043/0.043+0.000 0.127/0.138+0.010 PointNetVLAD [6] 0.841/0.845+0.004 0.442/0.451+0.009 0.663/0.738+0.075 0.934/0.978+0.044 0.123/0.234+0.111 0.541/0.600+0.060 PNPGAP(ours) 0.826/0.841+0.015 0.395/0.448+0.053 0.683/0.659-0.024 0.938/0.971+0.033 0.248/0.308+0.060 0.566/0.597+0.031
IV-E3 Cross Domain
The results of the cross-domain experiment are presented in Table VI. In this experiment, the performance difference between PNPGAP and PointNetVLAD is smaller, with PNPGAP outperforming by 3.3 pp. This difference is primarily attributable to PNPGAP’s superior performance on sequence 08, the most challenging sequence, where most revisits occur from the reverse direction. This finding suggests that, while the proposed aggregation approach provides less of an advantage in urban environments compared to horticultural settings, it still demonstrates superior rotation-invariant properties relative to other SOTA models, particularly PointNetVLAD.
When incorporating the SLC model during training, the performance improvements observed across models are consistent with those seen within the same domain, with the exception of OverlapTransformer. Notably, OverlapTransformer’s performance is significantly higher in the cross-domain experiment compared to its results on the HORTO-3DLM dataset, despite being trained on the same dataset in both cases. We hypothesize that this performance gain is primarily due to OverlapTransformer’s input representation, which appears to be better suited to the structured nature of urban environments.
@@ IV-F Runtime Analysis
Processing capacity and model size are critical factors for real-world implementation, particularly in resource-constrained robotics applications. The runtime experiment results, presented in Table VII, are obtained using the hardware detailed in Section IV-D. The reported times, measured in milliseconds (ms), represent the duration required by each model to process a batch of 20 scans. Retrieval duration was excluded from this experiment, as it remains consistent across all models. Model size is expressed by the number of parameters, in millions (M). PNPGAP is notably compact, with 0.4 million parameters, and processes a batch in an average of 4.4 ms.
Comparative analysis with other SOTA models reveals that PNPGAP is approximately 5% the size of LOGG3D-Net and 31 times faster. Compared to PointNetVLAD, PNPGAP is only 2% of its size and 6 times faster. These results underscore the computational efficiency of the proposed PNPGAP model.
TABLE VII: The average runtime per batch, where the batch size is 20 scans.
@@ V CONCLUSIONS
In this letter, we addressed the challenge of 3D LiDAR-based place recognition in horticultural environments by introducing the PointNetPGAP model, which integrates two statistically-inspired aggregators, along with an SLC model designed to mitigate inter-row descriptor ambiguities. We also presented the HORTO-3DLM dataset, comprising sequences from orchards and a strawberry plantation.
Experimental results obtained from the HORTO-3DLM dataset, using a cross-validation protocol, as well as from the KITTI Odometry dataset, suggest that PointNetPGAP outperforms SOTA models, as evidenced by an ablation study and benchmarks at both the sequence and segment levels. Furthermore, incorporating the SLC model during training consistently improved performance, particularly in segments with high inter-row ambiguity, highlighting the critical role of context-specific training signals.